nancyfelipe1:

After 1 year, a bank account contains $7,746.90. After 2 years, the same account contains $7,901.84. Assuming no deposits or withdrawals are made, which equation can be used to find y, the amount of money in the account after x years? (Round money values to the nearest penny.) y = 7591.96(1.02)x y = 7591.96(0.02)x y = 7595(1.02)x y = 7595(0.02)x

NonportritV2:

y = 7595(1.02)x y = 7,595(1.02)¹ = 7,746.90 y = 7,595(1.02)² = 7,901.84 Therefore the answer is option C "y = 7595(1.02)x." Good luck!